# Factor Augmented Tensor-on-Tensor Neural Networks

### v1

@@ 4 Numerical Results
4.1 Simulation Studies
We carry out simulation studies under different scenarios with various configurations of sample size, data dimensions, tensor ranks, and relationships between covariates and responses. Recall that ${\cal X}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$, ${\cal Y}_{t}\in\mathbb{R}^{p_{1}\times\cdots\times p_{q}}$, for $1\leq t\leq n$.
Data Generating Process. We begin with generating the low-rank core tensor ${\cal F}_{t}$ from a tensor autoregressive model ${\rm vec}({\cal F}_{t})=\Phi\cdot{\rm vec}({\cal F}_{t-1})+{\rm vec}({\cal W}_{t})$ for $t=1,\ldots,n$. Here, ${\rm vec}({\cal F}_{t})$ denotes the vectorization [^4] We would like to clarify that our proposed FATTNN approach does not involve tensor vectorization in the model fitting. This vectorization is solely for the convenience of data generation in simulation. of a tensor ${\cal F}_{t}$, ${\cal W}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is an error tensor in which every element is randomly generated from a normal distribution. The coefficient matrix $\Phi\in\mathbb{R}^{(r_{1}r_{2}\cdots r_{K})\times(r_{1}r_{2}\cdots r_{K})}$ is constructed using the Kronecker product of a sequence of matrices $Q_{k}$, i.e., $\Phi=\odot_{k=1}^{K}Q_{k}$, in which each $Q_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$ is a matrix with i.i.d. standard normal entries and orthonormalized through QR decomposition. The tensor time series ${\cal F}_{t}$ is randomly initiated in a distance past, and we discard the first 500 time points to stabilize the process.
@@ 4.1 Simulation Studies
We consider the following three experimental configurations:
  1. (1)
    $(d_{1},d_{2},d_{3})=(25,25,12)$,$(r_{1},r_{2},r_{3})=(6,8,6)$, $(p_{1},p_{2},p_{3})=(3,3,2)$
    $n=500$, $s(z)=cos(z)$, $\sigma_{u}^{2}=1$.
  2. (2)
    $(d_{1},d_{2},d_{3})=(30,6,12)$, $(r_{1},r_{2},r_{3})=(6,3,2)$, $(p_{1},p_{2},p_{3})=(8,6,4)$,
    $n=400$, $s(z)=log(|z|)$, $\sigma_{u}^{2}=1$.
  3. (3)
    $(d_{1},d_{2},d_{3})=(12,3,12)$, $(r_{1},r_{2},r_{3})=(4,3,4)$, $(p_{1},p_{2},p_{3})=(3,3,3)$,
    $n=100$, $s(z)=log(e^{(z)}+1)$, $\sigma_{u}^{2}=0.5$.
@@
Simulation Evaluations. In each setting, we use 70% for training the models and 30% for model evaluation. Comparisons of MSE, and computational time are summarized in Table 1 and Table 2. All reported metrics are averaged over 20 replications. Experiments were run on the research computing cluster provided by the University Center for Research Computing. From Table 1, we can see that FATTNN improves over the two benchmark methods by a large margin. The improvements can be attributed to two facets: (a) the utilization of tensor factors, summarizing useful predictive information effectively and reducing the number of noise covariates, and (b) the utilization of neural networks, possessing extraordinary ability in capturing complex relationships in the data and therefore exhibiting exceptional predictive power. On a separate note, Table 2 reveals the evident computational advantage of FATTNN compared to other neural network models such as TCN, owing to the effective dimension reduction through tensor factor models.
@@
Table 1: Comparisons of prediction accuracy of different methods in simulation studies
(measured by MSE over the testing data)
Prediction Task TCN FATTNN MultiWay
Simulation 1 3.642 (1.91) 2.451 (1.57) 3.987 (2.00)
Simulation 2 670.5 (25.8) 511.9 (22.5) 805.1 (28.4)
Simulation 3 1074 (32.6) 696.7 (26.4) 1670 (40.3)
@@
Note: The results are averaged over 20 replications.
Numbers in the brackets are standard errors.
Table 2: Comparisons of computational time (in the format of hh:mm:ss)
of different methods in simulation studies
Prediction Task TCN FATTNN Muiltway
Simulation 1 01:22:02 00:16:18 00:02:15
Simulation 2 01:26:30 00:17:54 00:02:30
Simulation 3 00:11:16 00:03:41 00:01:03
4.2 Real Data Examples
@@
We evaluate the performance of different methods using five prediction tasks on three real-world datasets. The information about sample sizes in the datasets, data dimensions of covariates $(d_{1},\ldots,d_{K})$ and responses $(p_{1},\ldots,p_{q})$, and the ranks of core tensors $(r_{1},\ldots,r_{K})$ utilized in the analyses are summarized in Table 3. For the prediction tasks using the New York Taxi data and FMRI image datasets, we split the data into training sets (70%) and testing (30%). For the prediction tasks using the FAO dataset, we use 80% data for training and 20% for testing due to its relatively small sample size. We compare the prediction accuracy and computation time of three methods: FATTNN, TCN, and Multiway. The numerical prediction results along with the corresponding computation time are summarized in Table 4.
@@
NYC taxi – midtown 504 (2,12,12,8) (2,12,12,8) (2,4,4,2)
NYC taxi – downtown 504 (2,19,19,8) (2,19,19,8) (2,4,4,2)
@@
Table 4: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time (in the format of hh:mm:ss) of different methods in real-world applications
MSE Time
Prediction Task TCN FATTNN Multiway TCN FATTNN Multiway
FAO crop $\sim$ crop 14.11 7.437 12.29 00:55:33 00:15:28 00:03:18
FAO $\sim$ livestock 32.53 21.65 48.04 00:24:34 00:12:57 00:03:21
NYC taxi – midtown 12.87 7.399 22.84 01:57:15 00:34:34 00:04:52
NYC taxi – downtown 14.58 9.493 28.10 03:25:55 00:32:46 00:25:35
FMRI data 0.1236 0.06634 0.1397 03:54:50 00:20:58 10:01:20
@@
(1) The United Nations Food and Agriculture Organization (FAO) Crops and Livestock Products Data. The database provides agricultural statistics (including crop, livestock, and forestry sub-sectors) collected from countries and territories since 1961. It is publicly available at https://www.fao.org/faostat/en/#data/QCL. Our analyses focus on 11 crops and 5 livestock products of 46 countries in East Asia, North America, South America, and Europe from 1961 to 2022. Detailed information about the types of crops, livestock, and countries are presented in Table S.1 of Appendix B. We study two prediction tasks using the FAO dataset: (i) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 selected countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$); and (ii) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$). Observing high variability in the raw data, we adopt a log transformation on the raw series and fit our models using log-transformed data.
Table 4 shows that FATTNN brings in substantial reductions in MSE compared to TCN and Multiway. In task (i), FATTNN yields 46.92% and 39.49% reductions in MSE compared to TCN and Multiway, respectively, and in task (ii), the improvements are 33.45% and 54.93% reductions in MSE, respectively. As reflected by the table, FATTNN computes much faster than TCN, about 3.6 times faster in task (i) and 2 times faster in tasks (ii). We also provide graphical comparisons between ground truth and predictions from different methods; see Figures 2 and 3. The plots confirm that FATTNN yields results that most closely align with the ground-truth patterns.
@@
Figure 2: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe. Values are reported on the log-transformed scale. From left to right: Ground truth, FATTNN, and TCN.
Figure 3: Comparisons between ground-truth values and predicted Yield (top row) and Area-harvested (bottom row) of citrus fruit in 26 selected countries in Europe using livestock data of the same 26 countries. Values are plotted on the log-transformed scale. For readability, numerical numbers are reported in Tables S.2S.3 instead of being superimposed on the plots. From left to right: Ground truth, FATTNN, and TCN.
@@
(2) New York City (NYC) Taxi Trip Data. The data contains 24-hour taxi pick-up and drop-off information of 69 areas in New York City for all the business days in 2014 and 2015. It is publicly available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page. Considering the autoregressive nature of this taxi data, we include the lag-1 response as an additional covariate when fitting the models. We focus on two prediction tasks: using the pick-up and drop-off data from 6:00-14:00 to predict the pick-up and drop-off outcomes from 14:00-22:00 in 12 districts in Midtown Manhattan and 19 districts in Downtown Manhattan, respectively. A detailed Manhattan district map is provided in Figure S.1 of Appendix B. The Midtown prediction task yields a tensor covariate and response $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 12\times 12\times 8}\times\mathbb{R}^{2\times 12\times 12\times 8}$ for each business day, in which the 1st dimension represents two different taxi companies, the 2nd and 3rd dimensions encode the pick-up and drop-off districts, and the 4th dimension represents the hour of day. Similarly, the Downtown prediction task involves $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 19\times 19\times 8}\times\mathbb{R}^{2\times 19\times 19\times 8}$ for each business day.
@@
From Table 4, we observe that FATTNN significantly outperforms the other approaches in terms of prediction accuracy, with a 42.5% reduction in MSE compared to TCN and a 67.6% reduction in MSE compared to Multiway in Midtown traffic prediction. In Downtown traffic prediction, FATTNN shows a 34.9% reduction in MSE compared to TCN and a 66.2% reduction in MSE compared to Multiway. As shown in the table, FATTNN is much more computationally efficient than TCN, about 6.4 times faster than TCN in Downtown prediction and 3.4 times faster in Midtown prediction. Additionally, Figure 4 visualizes the comparisons using a 5-day moving average of ground-truth values and predicted pick-up and drop-off volumes by day (where we sum the number of pick-up and drop-off across 8-hour testing periods). The FATTNN-predicted lines appear to be closer to the ground-truth lines than TCN-predicted lines in most plots.
(3) Functional Magnetic Resonance Imaging (FMRI) Data. We consider the Haxby dataset (Haxby et al.,, 2001), a well-known public dataset for research in brain imaging and cognitive neuroscience, which can be retrieved from the “NiLearn” Python library. The data contains slices of $64\times 64$ FMRI brain images collected from 1452 samples. We use the first 25 slices of each sample to predict the 26th slice of that sample, i.e., $(\mathcal{X}_{t},\mathcal{Y}_{t})\in\mathbb{R}^{25\times 64\times 64}\times\mathbb{R}^{1\times 64\times 64}$. Similar to the previous two datasets, FATTNN significantly outperforms the other methods in both prediction accuracy and computational efficiency. Compared to TCN, FATTNNyields a 46.3% reduction in MSE and computes 11 times faster. To further illustrate the algorithm performances, we visualize the predicted images in Figure 5. It is evident that FATTNN produces images that more accurately resemble the ground truth.
In summary, our proposed methods, especially the FATTNN, achieve substantial increases in prediction accuracy compared to benchmark methods – with $\sim 40\%$ reductions in MSE compared to TCN, and even larger reductions compared to Multiway Regression. At the same time, the computational costs of FATTNN methods are much lower compared to TCN which directly feeds tensor observations into a convolutional network, ranging from 2 to 11 times faster than TCN.
@@
Figure 4: Comparisons between ground-truth values and predicted pick-up and drop-off volumes using various methods. “District A to B” denotes the traffic volume that passengers were picked up in District A and dropped off in District B. The district numbering is assigned according to the Manhattan district map shown in Figure S.1. Left panel: Midtown Manhattan; Right panel: Downtown Manhattan.
Figure 5: Comparison between ground-truth values and predicted outputs for selected samples in FMRI dataset. From left to right: Ground truth, FATTNN, and TCN.
5 Summary & Discussion
@@
In this paper, we propose FATTNN, a hybrid method that integrates a tensor factor model into deep neural networks, for tensor-on-tensor time series forecasting. The key advantage of FATTNN is its ability to exploit and utilize the underlying tensor structure among the covariate tensors. Through a low-rank tensor factor representation, we preserve the tensor structures and, in the meantime, drastically reduce the data dimensionality, leading to enhanced prediction results and reduced computational costs. This approach unifies the strengths of both factor models and neural networks. Numerical studies confirm that our proposed methods successfully achieve substantial increases in prediction accuracy and significant reductions in computational time, compared to traditional linear tensor-on-tensor regression models (which often fail to capture the intricate nonlinearity in covariate-response relationships) and conventional deep learning approaches (that often operate in a black-box manner without intention to making use of the inherent structures among observations).
Limitations. In this work, we focus on tensor-on-tensor time series forecasting. The integration of TIPUP and TCN, both designed specifically for handling temporal data, makes our proposed approach a powerful tool for forecasting tensor series over time. When the data are not time series, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized for modeling non-temporal data. This limitation can be easily rectified by adopting proper types of tensor factor models and neural networks in the implementation of FATTNN, depending on the nature of the prediction tasks and data types involved. Please see the paragraph on generalizations of FATTNN.
@@
Generalizations. Though we presented our models for the purpose of tensor-on-tensor time series forecasting, our proposed FATTNN framework is not confined to temporal data but can accommodate a variety of data types. The main idea of FATTNN is to use a low-rank tensor factor model to capture the intrinsic patterns among the observed covariates, and then proceed with a neural network to model the intricate relationships between covariates and responses. With appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, graph, network, text, video). When handling time series tensors, we adopt TIPUP factorization and TCN to fully capitalize on their temporal nature. To extend beyond time series data, different types of factor models and neural network architectures should be deployed, depending on the nature of the data. For example, for graph data, Graph Neural Network is a more suitable alternative to TCN. For text data, Transformer models could substitute for the TCN in our FATTNN framework. Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architecture for tensor-on-tensor prediction using diverse data types.
Appendices
Appendix A Tensor Factor Models
A.1 Notation

### v2

@@ Introduction
Simulation Studies
Data Generating Process. We carry out simulation studies under different scenarios with various configurations of sample size, data dimensions, tensor ranks, and relationships between covariates and responses. Recall that ${\cal X}_{t}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$, ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$, ${\cal Y}_{t}\in\mathbb{R}^{p_{1}\times\cdots\times p_{q}}$, for $1\leq t\leq n$. We begin with generating the low-rank core tensor ${\cal F}_{t}$ from a tensor autoregressive model ${\rm vec}({\cal F}_{t})=\Phi\cdot{\rm vec}({\cal F}_{t-1})+{\rm vec}({\cal W}_{t})$ for $t=1,\ldots,n$. Here, ${\rm vec}({\cal F}_{t})$ denotes the vectorization [^3] We’d like to clarify that our proposed FATTNN approach does not involve tensor vectorization in the model fitting. This vectorization is solely for the convenience of data generation in simulation. of a tensor ${\cal F}_{t}$, ${\cal W}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is an error tensor in which every element is randomly generated from a normal distribution. The coefficient matrix $\Phi\in\mathbb{R}^{(r_{1}r_{2}\cdots r_{K})\times(r_{1}r_{2}\cdots r_{K})}$ is constructed using the Kronecker product of a sequence of matrices $Q_{k}$, i.e., $\Phi=\odot_{k=1}^{K}Q_{k}$, in which each $Q_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$ is a matrix with i.i.d. standard normal entries and orthonormalized through QR decomposition. The tensor time series ${\cal F}_{t}$ is randomly initiated in a distance past, and we discard the first 500 time points to stabilize the process.
@@
We consider the following experimental configurations:
(1) $(d_{1},d_{2},d_{3})=(25,25,12)$,$(r_{1},r_{2},r_{3})=(3,3,2)$, $(p_{1},p_{2},p_{3})=(6,8,6)$, $n=500$, $s(z)=cos(z)$, $\sigma_{u}^{2}=1$.
(2) $(d_{1},d_{2},d_{3})=(30,6,12)$, $(r_{1},r_{2},r_{3})=(6,3,2)$, $(p_{1},p_{2},p_{3})=(8,6,4)$, $n=400$, $s(z)=log(|z|)$, $\sigma_{u}^{2}=1$.
(3) $(d_{1},d_{2},d_{3})=(12,3,12)$, $(r_{1},r_{2},r_{3})=(4,3,4)$, $(p_{1},p_{2},p_{3})=(3,3,3)$, $n=100$, $s(z)=log(e^{(z)}+1)$, $\sigma_{u}^{2}=0.5$.
@@
Simulation Evaluations. In each setting, we use 70% for training the models and 30% for model evaluation. Comparisons of MSE, and computational time are summarized in Table 1. All reported metrics are averaged over 20 replications. Experiments were run on the research computing cluster provided by the University Center for Research Computing. From Table 1, we can see that FATTNN consistently improves over the rest benchmark methods by a large margin. The improvements can be attributed to two facets: (a) the utilization of tensor factors, summarizing useful predictive information effectively and reducing the number of noise covariates, and (b) the utilization of neural networks, possessing extraordinary ability in capturing complex relationships in the data and therefore exhibiting exceptional predictive power. On a separate note, Table 1 reveals the evident computational advantage of FATTNN compared to other neural network models such as TCN, owing to the effective dimension reduction through tensor factor models.
@@
Table 1: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time (in the format of hh:mm:ss) in simulation studies
Prediction Task Simulation 1 Simulation 2 Simulation 3
Test MSE
TCN 9.812 (3.13) 670.5 (25.8) 1074 (32.6)
FATTNN 6.353 (2.52) 511.9 (22.5) 696.7 (26.4)
MultiWay 12.31 (3.51) 805.1 (28.4) 1670 (40.3)
LSTM 10.48 (3.23) 878.3 (29.6) 807.8 (28.3)
TRL 12.29 (3.50) 852.6 (29.0) 2003 (38.4)
Conv-TT-LSTM 10.32 (3.15) 1643 (39.5) 2812 (52.8)
Computational Time
TCN 01:22:02 01:26:30 00:11:16
FATTNN 00:16:18 00:17:54 00:03:41
MultiWay 00:20:15 00:12:30 00:01:03
LSTM 00:03:36 00:03:01 00:02:09
TRL 00:02:25 00:01:30 00:00:50
Conv-TT-LSTM 03:02:15 05:31:25 00:35:37
@@
Note: The results are averaged over 20 replications. Numbers in the brackets are standard errors.
Real Data Examples
@@
We evaluate the performance of different methods using five prediction tasks on three real-world datasets. The information about sample sizes in the datasets, data dimensions of covariates $(d_{1},\ldots,d_{K})$ and responses $(p_{1},\ldots,p_{q})$, and the ranks of core tensors $(r_{1},\ldots,r_{K})$ utilized in the analyses are summarized in Table 2. For the prediction tasks using the New York Taxi data and FMRI image datasets, we split the data into training sets (70%) and testing (30%). For the prediction tasks using the FAO dataset, we use 80% data for training and 20% for testing due to its relatively small sample size. We compare the prediction accuracy and computation time of six methods illustrated in the simulation section. The numerical prediction results along with the corresponding computational time are summarized in Table 3.
@@
NYC taxi – midtown 504 (2,12,12,8) (12,12,8) (2,4,4,2)
NYC taxi – downtown 504 (2,19,19,8) (19,19,8) (2,4,4,2)
@@
Table 3: Comparisons of prediction accuracy (measured by MSE over the testing data) and computational time
(in the format of hh:mm:ss) of different methods in real-world applications
Prediction Task TCN FATTNN Multiway TRL LSTM Conv-TT-LSTM
Test MSE
FAO crop $\sim$ crop 14.11 [10.83, 17.58] 7.437 [6.18, 11.44] 12.29 [11.86, 32.45] 17.84 [15.89, 22.74] 15.76 [13.16, 18.34] 53.39 [52.41, 63.95]
FAO $\sim$ livestock 32.53 [24.87, 36.10] 21.65 [19.56, 25.10] 48.04 [45.56, 165.80] 66.25 [56.85, 76.13] 34.23 [30.47, 38.48] 69.61 [69.06, 78.44]
NYC taxi – midtown 12.87 [12.06, 15.52] 7.399 [6.53, 8.03] 22.84 [20.86, 44.79] 8.565 [7.89, 10.12] 8.404 [7.95, 9.63] 25.80 [21.55, 28.62]
NYC taxi – downtown 14.58 [13.36, 17.42] 9.493 [8.46, 11.03] 28.10 [23.99, 52.12] 12.43 [10.82, 14.50] 12.45 [11.09, 14.38] 42.18 [36.27, 47.66]
FMRI data 0.1236 [0.0997, 0.148] 0.06634 [0.053, 0.075] 0.1397 [0.135, 0.158] 0.07902 [0.072, 0.084] 0.1964 [0.163, 0.224] 0.03871 [0.0337, 0.0914]
Computational Time
FAO crop $\sim$ crop 00:55:33 00:15:28 00:03:18 00:09:41 00:02:08 00:20:07
FAO $\sim$ livestock 00:24:34 00:12:57 00:03:21 00:07:53 00:01:55 00:16:18
NYC taxi – midtown 01:57:15 00:34:34 00:04:52 00:19:26 00:05:01 01:54:28
NYC taxi – downtown 03:25:55 00:32:46 00:25:35 00:22:09 00:05:34 03:07:05
FMRI data 03:54:50 00:20:58 10:01:20 00:43:51 00:06:27 17:27:05
Note: Numbers in the square brackets are bootstrap confidence intervals (CIs) of test MSE over 100 bootstrapping replications.
@@
(1) The United Nations Food and Agriculture Organization (FAO) Crops and Livestock Products Data. The database provides agricultural statistics (including crop, livestock, and forestry sub-sectors) collected from countries and territories since 1961. It is publicly available at https://www.fao.org. Our analyses focus on 11 crops and 5 livestock products from 46 countries in East Asia, North America, South America, and Europe from 1961 to 2022. Detailed information about the types of crops, livestock, and countries are presented in Table S.1 of Appendix B. We study two prediction tasks: (i) using Yield and Production data of 11 different crops for 33 countries in East Asia, North America, and Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{33\times 2\times 11}$) to predict the Yield and Production quantities of the same crops for 13 selected countries in South America (i.e., ${\cal Y}_{t}\in\mathbb{R}^{13\times 2\times 11}$); and (ii) using four agricultural statistics (Producing Animals, Animals-slaughtered, Milk Animals, Laying) associated with 5 kinds of livestock of 26 selected countries in Europe (i.e., ${\cal X}_{t}\in\mathbb{R}^{26\times 4\times 5}$) to predict three metrics (Area-harvested, Production, and Yield) of 11 crops in the same countries (i.e., ${\cal Y}_{t}\in\mathbb{R}^{26\times 3\times 11}$). Observing high variability in the raw data, we adopt a log transformation on the raw series and fit our models using log-transformed data.
@@
Figure 2: Left Panel: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Yield of Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe; Right Panel: Comparisons between ground-truth values and predicted Area-harvested (top row) and Yield (bottom row) of citrus fruit in 26 selected countries in Europe using livestock data of the same 26 countries. Values are plotted on the log-transformed scale. For readability, numerical values from the figures are tabulated in Appendix B. In each panel, from left to right: Ground truth, FATTNN, and TCN.
Table 3 shows that FATTNN brings in substantial reductions in MSE compared to other methods. In task (i), FATTNN yields 47.3%, 39.5%, 58.3%, 52.8%, and 86.1% reductions in MSE compared to TCN, Multiway, TRL, LSTM, and Conv-TT-LSTM, respectively, and in task (ii), the improvements are 33.4%, 54.9%, 67.3%, 36.7%, and 68.9% reductions in MSE, respectively. The table also shows that the upper bound of the bootstrap CI for FATTNN is smaller than the lower bounds of the CIs for almost all the benchmarks, indicating FATTNN possesses a statistically significant increase in prediction accuracy. What’s more, FATTNN produces the shortest CI, implying the lowest variability among the methods compared. As reflected by the table, FATTNN computes much faster than TCN, about 3.6 times faster in task (i) and 2 times faster in task (ii). We also provide graphical comparisons between ground truth and predictions from different methods; see Figure 2. The plots confirm that FATTNN yields results that most closely align with the ground-truth patterns.
@@
(2) New York City (NYC) Taxi Trip Data. The data contains 24-hour taxi pick-up and drop-off information of 69 areas in New York City for all the business days in 2014 and 2015. It is publicly available at https://www.nyc.gov. Considering the autoregressive nature of this taxi data, we include the lag-1 response as an additional covariate when fitting the models. We focus on two prediction tasks: using the pick-up and drop-off data from 6:00-14:00 to predict the pick-up and drop-off outcomes from 14:00-22:00 in 12 districts in Midtown Manhattan and 19 districts in Downtown Manhattan, respectively. A detailed Manhattan district map is provided in Figure S.3 of Appendix B. The Midtown prediction task yields a tensor covariate and response $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 12\times 12\times 8}\times\mathbb{R}^{12\times 12\times 8}$ for each business day, in which the 1st dimension stands for the observed information at time $t$, concatenated with lag 1 response, the 2nd and 3rd dimensions encode the pick-up and drop-off districts, and the 4th dimension represents the hour of day. Similarly, the Downtown prediction task involves $({\cal X}_{t},{\cal Y}_{t})\in\mathbb{R}^{2\times 19\times 19\times 8}\times\mathbb{R}^{19\times 19\times 8}$for each business day.
@@
From Table 3, we observe that FATTNN significantly outperforms all other approaches in terms of prediction accuracy, with 42.5%, 67.6%, 13.6%, 12.0%, and 71.3% reductions in MSE compared to TCN, Multiway, TRL, LSTM, and Conv-TT-LSTM, respectively, in Midtown traffic prediction. In Downtown traffic prediction, FATTNN shows 34.9%, 66.2%, 23.6%, 23.7%, and 77.5% reductions in MSE compared to TCN, Multiway, TRL, LSTM, and Conv-TT-LSTM, respectively. As shown in the table, FATTNN is much more computationally efficient than TCN which directly feeds tensor observations into a convolutional network, about 6.4 times faster than TCN in Downtown prediction and 3.4 times faster in Midtown prediction. Additionally, Figure 3 visualizes the comparisons using a 5-day moving average of ground-truth values and predicted pick-up and drop-off volumes by day (where we sum the number of pick-up and drop-off across 8-hour testing periods). The FATTNN-predicted lines appear to be closer to the ground-truth lines than TCN-predicted lines in most plots.
@@
Figure 3: Comparisons between ground-truth values and predicted pick-up and drop-off volumes using various methods. “District A to B” denotes the traffic volume that passengers were picked up in District A and dropped off in District B. The district numbering is assigned according to the Manhattan district map shown in Figure S.3. An enlarged figure and more detailed descriptions are in Appendix B. Left: Midtown Manhattan; Right: Downtown Manhattan.
(3) Functional Magnetic Resonance Imaging (FMRI) Data. We consider the Haxby dataset (Haxby et al. 2001), a well-known public dataset for research in brain imaging and cognitive neuroscience, which can be retrieved from the “NiLearn” Python library. The data contains slices of $64\times 64$ FMRI brain images collected from 1452 samples. We use the first 25 slices of each sample to predict the 26th slice of that sample, i.e., $(\mathcal{X}_{t},\mathcal{Y}_{t})\in\mathbb{R}^{25\times 64\times 64}\times\mathbb{R}^{1\times 64\times 64}$.
Conv-TT-LSTM yields the lowest MSE in this task. This is not surprising as Conv-TT-LSTM is specifically designed for image prediction while FATTNN is more suitable for time-series tensor prediction. Other than Conv-TT-LSTM, FATTNN still outperforms other methods. Additionally, though Conv-TT-LSTM has the best prediction accuracy, it is quite computationally heavy - about 50 times slower than FATTNN. The FATTNN is also 2 times faster than the TRL method which has the 3rd lowest MSE in this task.
In summary, our proposed FATTNN performs the best in the four prediction tasks on FAO and NYC datasets. In the FMRI dataset, the FATTNN yields higher MSE than Conv-TT-LSTM, a method specialized in image prediction. That being said, FATTNN has a huge computational advantage over Conv-TT-LSTM. We also observe that Conv-TT-LSTM underperforms other methods in the FAO and NYC datasets by a large margin, when the prediction tasks are not focusing on image data. To summarize, the FATTNN demonstrates the overall best performance on all the prediction tasks, and strikes an ideal balance between prediction accuracy and computational cost. The FATTNN achieves substantial increases in prediction accuracy compared to benchmark methods. At the same time, the computational costs of FATTNN methods are much lower compared to TCN and Conv-TT-LSTM, ranging from 2 to 50 times faster.
Summary & Discussion
@@
In this paper, we propose FATTNN, a hybrid method that integrates a tensor factor model into deep neural networks, for tensor-on-tensor time series forecasting. The key advantage of FATTNN is its ability to exploit and utilize the underlying tensor structure among the covariate tensors. Through a low-rank tensor factor representation, we preserve the tensor structures and, in the meantime, drastically reduce the data dimensionality, leading to enhanced prediction results and reduced computational costs. This approach unifies the strengths of both factor models and neural networks. Numerical studies confirm that our proposed methods successfully achieve substantial increases in prediction accuracy and significant reductions in computational time, compared to traditional linear tensor-on-tensor regression models (which often fail to capture the intricate nonlinearity in covariate-response relationships), conventional deep learning methods (that often operate in a black-box manner without making use of the inherent structures among observations), and state-of-the-art tensor-decomposition-based learning methods.
@@
Limitation and Extension. Though we presented our models for the purpose of tensor-on-tensor time series forecasting, our proposed FATTNN framework is not confined to temporal data but can accommodate a variety of data types. The main idea of FATTNN is to use a low-rank tensor factor model to capture the intrinsic patterns among the observed covariates, and then proceed with a neural network to model the intricate relationships between covariates and responses. When handling time series tensors, we adopt TIPUP factorization and TCN to fully capitalize on their temporal nature. For non-temporal data, the TIPUP-TCN-based FATTNN is still applicable for tensor-on-tensor prediction tasks, but the strengths of TIPUP and TCN may not be fully utilized. For higher-order tensors (e.g. $K>4$ ), Tensor-Train (Oseledets 2011) and Hierarchical Tucker decompositions (Lubich et al. 2013) may be better suited. This limitation can be easily rectified: with appropriate choices of factor models and neural network architectures, the proposed FATTNN can be naturally adapted to any simple (e.g., i.i.d observations) or complex tensor-type data (e.g., image, network, text, video). To extend beyond time series data, different types of factor models and neural network architectures should be deployed, depending on the nature of the data. For example, for graph data, Graph Neural Network is a more suitable alternative to TCN. For text data, Transformer models could substitute for the TCN in our FATTNN framework. An extended discussion on the generality of the FATTNN framework is presented in Appendix C. Our proposed FATTNN is flexible in accommodating different types of tensor factor models and deep learning architectures for tensor-on-tensor prediction using diverse data types.
Acknowledgments
We thank the anonymous reviewers for helpful suggestions. The work of Yuefeng Han is supported in part by National Science Foundation Grants DMS-2412578. The work of Xiufan Yu is supported in part by National Institutes of Health grant R01GM152812.
Appendix A Appendix A: Tensor Factor Models
Notation